a shopkeeper buys a number of books for rs 800 .if he had bought 4 more books for the same amount each book would ave cost him 10 rs less.how many books did he buy?

Grade:12

1 Answers

Sharan
9 years ago
let cost of each book=x
let no of books bought be = y
Therefore according to question,,
xy=800.................(1)
and,   (x-10)(y+4)=800..............(2)
xy+4x-10y-40=800
800+4x-10y-40=800
4x-10y-40=0.............(3)
Now using (1)&(3) a quadratic equation will form and solve it for x&y